# Shaocheng Liu and Qi Chen

### v2

@@ II-C Entropy functions on the extreme rays of $\Gamma_{4}$
$X_{3}\mod k=(X_{2}+\lfloor\dfrac{X_{1}}{k}\rfloor)\mod k.$ (II.C.27)
Note that $X_{3}$ is distributed on $\{0,1,\dots,k-1\}$, we have
$X_{3}=(X_{2}+\lfloor\dfrac{X_{1}}{k}\rfloor)\mod k,$ (II.C.28)
@@
$X_{2}=(X_{3}-\lfloor\dfrac{X_{1}}{k}\rfloor)\mod k.$ (II.C.29)
By (II.C.26), replacing $(X_{3}-X_{2})\mod k$ by $\lfloor\dfrac{X_{1}}{k}\rfloor$ in (II.C.23),
$X_{1}=k\lfloor\dfrac{X_{1}}{k}\rfloor+[(X_{4}-X_{2})\mod k].$ (II.C.30)
@@
$X_{4}$ $=(X_{1}+X_{2}-\lfloor\dfrac{X_{1}}{k}\rfloor)\mod k,$ (II.C.31)
$X_{2}$ $=(X_{4}-X_{1}+\lfloor\dfrac{X_{1}}{k}\rfloor)\mod k$ (II.C.32)
Based on (II.C.28), (II.C.29), (II.C.31) and (II.C.32), the random vetor $(X_{i},i\in N_{4})$ must satisfy the following equalities,
$H(X_{123})$ $=H(X_{12})=H(X_{13}),$ (II.C.33)
$H(X_{124})$ $=H(X_{12})=H(X_{14}).$ (II.C.34)
@@
$H(X_{1234})=H(X_{234}),$ (II.C.38)
and (II.C.25), we have
$H(X_{12})=3\log k.$ (II.C.39)
By (II.C.37) and (II.C.33), $3\log k=H(X_{13}\leq H(X_{134})\leq H(X_{1234})=3\log k$ which implies
$H(X_{134})=3\log k.$ (II.C.40)
@@
$H(X_{23})=H(X_{24})=H(X_{34})=2\log k$ (II.C.41)
By (II.C.24), (II.C.39), (II.C.33), (II.C.34), (II.C.41), (II.C.37), (II.C.38) and (II.C.40) the value of $H(X_{A})$ are determined for all $A\subseteq N$ except for $A=\{1\}$. Now we verify that $H(X_{1})=2\log k$. By (II.C.24) and (II.C.39),
$H(X_{1})\geq H(X_{12})-H(X_{2})=2\log k.$ (II.C.42)
On the other hand, as $X_{1}$ is distributed on $\{0,1,\dots,k^{2}-1\}$ by (II.C.23), we have $H(X_{1})\leq 2\log k$. The proof is accomplished. ∎
@@ III Two-dimensional faces of $\Gamma_{4}$
6:         if  the $k$-th facet $F_{k}$ contains both $E_{i}$ and $E_{j}$then
@@
The result of Algorithm 1, i.e., $2$-dimensional faces corresponding to those “$1$”-entries in $C$ are classified in Table 2. Similar to the fact that the $41$ extreme rays can be classified in $11$ types, the $2$-dimensional faces can also be classified to $59$ types, and in each type, each face can be obtained from another by permuting the indices in $N_{4}$. These $59$ types of faces are listed in the upper triangle of Table 2. We label the rows and columns with the $11$ types extreme rays of $\Gamma_{4}$. For simplicity, we denote the face $F=\mathrm{cone}(E_{1},E_{2})$ by $(E_{1},E_{2})$. In each cell with “$(E_{1},E_{2})\ n$, Thm. $m$”, “$(E_{1},E_{2})$” denote a representative of the type of $2$-dimensional faces, where “$E_{1}$ $(E_{2})$” is a representative of the type the extrem rays in the column (row), “$n$” is the number of the faces in this type and this face is characterized in “Thm. $m$” in Section IV. For the cell with “00”, the convex hull of the two extreme rays in each type forms no $2$-dimensional faces of $\Gamma_{4}$. More details of the faces in Table 2 are discussed in Section IV.
@@
$({U^{1}_{1,1},U^{2}_{1,1})}$ $12$,Thm.IV-A.1 $12$, Thm.IV-A.1 $({U_{1,4},U^{1}_{1,1})}$ $12$, Thm.IV-B.2 $12$, Thm.IV-B.2 $(U_{2,4},U^{1}_{1,1})$ $(U_{3,4},U^{1}_{1,1})$ $4$,Thm.IV-B.3 $4$,Thm.IV-B.5 $12$,Thm.IV-G.1
$6$, Thm.IV-A.1 $({U^{12}_{1,2},U^{3}_{1,1})}$ $({U^{123}_{1,3},U^{4}_{1,1})}$ $4$, Thm.IV-A.1 $(U_{2,3}^{123},U_{1,1}^{4})$ $(\mathcal{W}_{2}^{34},U_{1,1}^{1})$ $4$ $4$, Thm.IV-E.1 $(\hat{U}^{1}_{2,5},U_{1,1}^{2})$ $(\hat{U}^{1}_{3,5},U^{2}_{1,1})$ $(V^{12}_{8},U^{3}_{1,1})$
$12$, Thm.IV-A.1 $4$, Thm.IV-A.1 $4$, Thm.IV-B.2 $12$, Thm.IV-B.2 $12$, Thm.IV-B.3 $12$, Thm.IV-B.5 $12$, Thm.IV-G.1
@@
$(U^{12}_{1,2},U^{13}_{1,2})$ $({U^{123}_{1,3},U^{12}_{1,2})}$ $(U_{2,3}^{123},U_{1,2}^{12})$ $6$, Thm.IV-B.2
$(U^{12}_{1,2},U^{34}_{1,2})$ $({U^{123}_{1,3},U^{14}_{1,2})}$ $6$, Thm.IV-A.1 $(U_{2,3}^{123},U_{1,2}^{14})$ $24$, Thm.IV-B.2 $6$ $6$, Thm.IV-E.2 $12$, Thm.IV-B.3 $12$, Thm.IV-B.5 $24$, Thm.IV-G.1
$3$, Thm.IV-A.1 $12$, Thm.IV-A.1 $12$, Thm.IV-B.2 $(\mathcal{W}_{2}^{34},U_{1,2}^{12})$
$6$, Thm.IV-B.1
[…]
$6$, Thm.IV-A.1 $4$, Thm.IV-A.1 $12$, Thm.IV-C.1 $12$, Thm.IV-B.1 $4$ $4$, Thm.IV-E.4 $12$, Thm.IV-B.4 $4$, Thm.IV-F.1 $12$, Thm.IV-G.1
[…]
$4$, Thm.IV-C.3 $1$, Thm.IV-E.5 $6$, Thm.IV-G.1
[…]
$6$, Thm.IV-C.2 $12$, Thm.IV-D.1 $4$ $4$, Thm.IV-E.3 $4$ $12$, Thm.IV-F.2 $12$, Thm.IV-G.1
[…]
$12$ $6$ $12$ $12$, Thm.IV-F.3
[…]
$4$ $4$
[…]
$6$, Thm.IV-G.1
[…]
@@ IV Characterization of entropy functions on two-dimensional faces of $\Gamma_{4}$
In this section, we characterize the entropy functions on $2$-dimensional faces of $\Gamma_{4}$. We embed each face $F=(E_{1},E_{2})$ in the first octant of a $2$-dimensional cartesian coordinate system whose axis are labeled by $a$ and $b$. Thus, for each $(a,b)$, $a,b\geq 0$, it represents the polymatroid $a\mathbf{r}_{1}+b\mathbf{r}_{2}$, where $\mathbf{r}_{i},i=1,2$ is the rank function of the minimal integer polymatroid in $E_{i}$, respectively. Throughtout this paper, for a random vector $(X_{i},i\in N_{4})$ or its subvectors, we assume each $X_{i}$ is distributed on a finite set $\mathcal{X}_{i}$, and for each $x_{i}\in X_{i}$, $p(x_{i})>0$.
@@
For $(x_{i},i\in N_{3})\in\mathcal{X}_{N_{3}}$ with $p(x_{123})>0$, the information equalties imply that the probability mass function satisfies
@@
Now we prove that each connected component is a complete tripartite graph. In a connected component, for any two vertices $x_{i}\in\mathcal{X}_{i},x_{j}\in\mathcal{X}_{j}$, $i,j\in N_{3},i\neq j$, there exists a path from $x_{i}$ to $x_{j}$ and the length of the path is denoted by $n$. We prove that $x_{i}$ and $x_{j}$ are adjacent by induction on $n$. First, it is obviously ture for $n=1$ and it holds for $n=2$ by Claim 1. Assume it is ture for some positive integers $k-1$ and $k$ where $k\geq 2$. Now consider $n=k+1$. If the vertex $y$ adjacent to $x_{j}$ on the path is not in $\mathcal{X}_{i}$, by induction hypothesis on $k$, we see that $x_{i}$ is adjacent to $y$. Then by Lemma 1, $x_{i}$ is adjacent to $x_{j}$. If $y$ is in $\mathcal{X}_{i}$, the vertex $z(z\neq x_{j})$ adjacent to $y$ on the path is adjacent to $x_{i}$ by induction on $k-1$. Then by Claim 2, $x_{i}$ is adjacent to $x_{j}$.
@@ IV-A All-entropic faces
Theorem IV-A.1 is also a corollary of [27, Proposition 4] or [28, Theorem 2].
@@ IV-B Entropy functions on the faces extended from $2$-dimensional faces of $\Gamma_{3}$
For $F=(U_{2,3}^{123},U_{1,2}^{12}),(\mathcal{W}_{2}^{34},U_{1,2}^{12})$, or $(\mathcal{W}_{2}^{14},U_{1,3}^{124})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a+b\geq\log\lceil 2^{a}\rceil$. [^3] We take the base of the logarithm be $2$ for computing the entropy.
@@
Let $F$ be any of the $3$ faces and $\mathbf{h}\in F$. Let $\mathbf{h}=a\mathbf{r}_{1}+b\mathbf{r}_{2}$, where $\mathbf{r}_{1}$ and $\mathbf{r}_{2}$ are the rank functions of the matroids on the two extreme rays of the face, respectively. Restricting $\mathbf{h}$ on $N_{3}$, we obtain $\mathbf{h}^{\prime}=a\mathbf{r}^{\prime}_{1}+b\mathbf{r}^{\prime}_{2}$, where $\mathbf{r}^{\prime}_{i},i=1,2$ are the restriction of $\mathbf{r}_{i}$ on $N_{3}$. It can be checked that they are the rank functions of $U_{2,3}$ and $U^{12,3}_{1,2}$, respectively. Thus, $\mathbf{h}^{\prime}\in(U_{2,3},U^{12,3}_{1,2})$ and $a+b\geq\log\lceil 2^{a}\rceil$ by [9, Theorem 1].
@@
For $F=(U_{2,3}^{123},U_{1,1}^{1}),(U_{2,3}^{123},U_{1,1}^{4}),(U_{2,3}^{123},U_{1,2}^{14}),$ $(\mathcal{W}_{2}^{14},U_{1,1}^{1}),(\mathcal{W}_{2}^{34},U_{1,1}^{1}),(\mathcal{W}_{2}^{14},U_{1,2}^{14})$ and $(\mathcal{W}_{2}^{24},U_{1,2}^{14})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a=\log{k}$ for integer $k>0$.
$a$$b$Olog2log3log4
Figure 2: The faces $(U_{2,3}^{123},U_{1,1}^{1}),(U_{2,3}^{123},U_{1,1}^{4}),(U_{2,3}^{123},U_{1,2}^{14}),$ $(\mathcal{W}_{2}^{14},U_{1,1}^{1}),(\mathcal{W}_{2}^{34},U_{1,1}^{1}),(\mathcal{W}_{2}^{14},U_{1,2}^{14})$ and $(\mathcal{W}_{2}^{24},U_{1,2}^{14})$
@@
For $F=(U_{2,3}^{123},U_{1,1}^{4})$ and $\mathbf{h}\in F$, restricting $\mathbf{h}$ on $N_{3}$, we obtain $\mathbf{h}^{\prime}=a\mathbf{r}^{\prime}_{1}$, where $\mathbf{r}^{\prime}_{1}$ is the rank function of $U_{2,3}$, the matroid on the extreme ray of $\Gamma_{3}$. Thus, if $\mathbf{h}$ is entropic, then $a=\log{k}$ for some integer $k>0$. Now for any $\mathbf{h}\in F$ with $a=\log{k}$, let $\mathbf{h}^{\prime}$ be its restriction on $N_{3}$ and $(X_{1},X_{2},X_{3})$ be the characterizing random variable of $\mathbf{h}^{\prime}$. Let $X_{4}$ be an arbitrary random variable independent of $(X_{1},X_{2},X_{3})$ with $H(X_{4})=b$. Then it can be checked that $(X_{i},i\in N_{4})$ is the characterizing random variable of $\mathbf{h}$. Thus $\mathbf{h}$ is entropic.
For $F$ be any of the other $6$ faces and $\mathbf{h}\in F$, let $\mathbf{h}=a\mathbf{r}_{1}+b\mathbf{r}_{2}$, where $\mathbf{r}_{1}$ and $\mathbf{r}_{2}$ are the rank functions of the matroids on the two extreme rays of the face, respectively. Restricting $\mathbf{h}$ on $N_{3}$, we obtain $\mathbf{h}^{\prime}=a\mathbf{r}^{\prime}_{1}+b\mathbf{r}^{\prime}_{2}$, where $\mathbf{r}^{\prime}_{i}=1,2$ is the restriction of $\mathbf{r}_{i}$ on $N_{3}$, and the rank function of $U_{2,3}$ and $U^{1,3}_{1,1}$, respectively. Thus, $\mathbf{h}^{\prime}\in(U_{2,3},U^{1,3}_{1,1})$ and $a=\log{k}$ for positive $k\in\mathbb{Z}$ by [10, Theorem 1].
Now we prove the “if ” part. Let $\mathbf{h}=a\mathbf{r}_{1}+b\mathbf{r}_{2}\in F$ and $a=\log{k}$ for some integer $k\geq 0$. Let $\mathbf{h}^{\prime}$ be the restriction of $\mathbf{h}$ on $N_{3}$. Therefore, $\mathbf{h}^{\prime}\in(U_{2,3},U^{1,3}_{1,1})$. Note that $\mathbf{h}^{\prime}=a\mathbf{r}_{1}+b\mathbf{r}_{2}$, where $\mathbf{r}_{1}$ and $\mathbf{r}_{2}$ are respectively the rank functions of $U_{2,3}$ and $U^{1,3}_{1,2}$. Let $(X^{\prime}_{1},X^{\prime}_{2},X^{\prime}_{3})$ be the characterizing random vector of $a\mathbf{r}_{1}$, that is, $X^{\prime}_{1}$ and $X^{\prime}_{2}$ are independent and both uniformly distributed on $N_{k}$, and $X^{\prime}_{3}=X^{\prime}_{1}+X^{\prime}_{2}\mod k$. Let $(X^{\prime\prime}_{1},X^{\prime\prime}_{2},X^{\prime\prime}_{3})$ be characterizing random vector of $b\mathbf{r}_{2}$, that is, $X^{\prime\prime}_{1}$ is an arbitrary random variable with $H(X^{\prime\prime}_{1})=b$, and $X^{\prime\prime}_{2}$ and $X^{\prime\prime}_{3}$ are both constant. Let $X_{i}=(X^{\prime}_{i},X^{\prime\prime}_{i})$, $i\in N_{3}$. Then it can be checked that $(X_{1},X_{2},X_{3})$ is the characterizing random vector of $\mathbf{h}^{\prime}$.
@@
For $F=(\hat{U}_{2,5}^{1},U_{1,1}^{1})$, $(\hat{U}_{2,5}^{1},U_{1,1}^{2})$ and $(\hat{U}_{2,5}^{1},U_{1,2}^{12})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a=\log k$ for integer $k>0$.

### v3

@@ II-C Entropy functions on the extreme rays of $\Gamma_{4}$
$X_{3}\mod v=(X_{2}+\lfloor\dfrac{X_{1}}{v}\rfloor)\mod v.$ (II.C.28)
Note that $X_{3}$ is distributed on $\mathbb{I}_{v}$, we have
$X_{3}=(X_{2}+\lfloor\dfrac{X_{1}}{v}\rfloor)\mod v,$ (II.C.29)
@@
$X_{2}=(X_{3}-\lfloor\dfrac{X_{1}}{v}\rfloor)\mod v.$ (II.C.30)
By (II.C.27), replacing $(X_{3}-X_{2})\mod v$ by $\lfloor\dfrac{X_{1}}{v}\rfloor$ in (II.C.23),
$X_{1}=v\lfloor\dfrac{X_{1}}{v}\rfloor+[(X_{4}-X_{2})\mod v].$ (II.C.31)
@@
$X_{4}$ $=(X_{1}+X_{2}-\lfloor\dfrac{X_{1}}{v}\rfloor)\mod v,$ (II.C.32)
$X_{2}$ $=(X_{4}-X_{1}+\lfloor\dfrac{X_{1}}{v}\rfloor)\mod v$ (II.C.33)
Based on (II.C.29), (II.C.30), (II.C.32) and (II.C.33), the random vetor $(X_{i},i\in N_{4})$ must satisfy the following equalities,
$H(X_{123})$ $=H(X_{12})=H(X_{13}),$ (II.C.34)
$H(X_{124})$ $=H(X_{12})=H(X_{14}).$ (II.C.35)
@@
$H(X_{1234})=H(X_{234}),$ (II.C.40)
and (II.C.26), we have
$H(X_{12})=3\log v.$ (II.C.41)
By (II.C.39) and (II.C.34), $3\log v=$ $H(X_{13})$$\leq H(X_{134})\leq H(X_{1234})=3\log v$ which implies
$H(X_{134})=3\log v.$ (II.C.42)
@@
$H(X_{23})=H(X_{24})=H(X_{34})=2\log v$ (II.C.43)
By (II.C.24), (II.C.41), (II.C.34), (II.C.35), (II.C.43), (II.C.39), (II.C.40) and (II.C.42) the value of $H(X_{A})$ are determined for all $A\subseteq N$ except for $A=\{1\}$. Now we verify that $H(X_{1})=2\log v$. By (II.C.24) and (II.C.41),
$H(X_{1})\geq H(X_{12})-H(X_{2})=2\log v.$ (II.C.44)
On the other hand, as $X_{1}$ is distributed on $\mathbb{I}_{v^{2}}$ by (II.C.23), we have $H(X_{1})\leq 2\log v$. The proof is accomplished. ∎
@@ III Two-dimensional faces of $\Gamma_{4}$
6:   if the $k$-th facet $F_{k}$ contains both $E_{i}$ and $E_{j}$, then
@@
The result of Algorithm 1, i.e., $510$ 22-dimensional faces corresponding to those “11”-entries in $C$ are classified in Table 3. Similar to the fact that the $41$ extreme rays can be classified in $11$ types, the $510$ 22-dimensional faces can also be classified to $59$ types, and in each type, each face can be obtained from another by permuting the indices in $N_{4}$. These $59$ types of faces are listed in the upper triangle of Table 3. We label the rows and columns with the $11$ types of extreme rays of $\Gamma_{4}$. For simplicity, we denote the face $F=\mathrm{cone}(E_{1},E_{2})$ by $(E_{1},E_{2})$. In each cell with $(E_{1},E_{2})\ n$, Thm. mm”, $(E_{1},E_{2})$” denote a representative of the type of $2$-dimensional faces, where $E_{1}$ $(E_{2})$” is a representative of the type the extrem rays in the column (row), “nn” is the number of the faces in this type and this face is characterized in “Thm. mm” in Section IV. For the cell with “0”, the convex hull of the two extreme rays in each type forms no $2$-dimensional faces of $\Gamma_{4}$. More details of the faces in Table 3 are discussed in Section IV.
@@
$({U^{1}_{1,1},U^{2}_{1,1})}$ $12$,Thm.IV-A.1 $12$, Thm.IV-A.1 $({U_{1,4},U^{1}_{1,1})}$ $12$, Thm.IV-B.2 $12$, Thm.IV-B.2 $(U_{2,4},U^{1}_{1,1})$ $(U_{3,4},U^{1}_{1,1})$ 44,Thm.IV-B.3 44,Thm.IV-B.5 $12$,Thm.IV-G.1
66, Thm.IV-A.1 $({U^{12}_{1,2},U^{3}_{1,1})}$ $({U^{123}_{1,3},U^{4}_{1,1})}$ 44, Thm.IV-A.1 $(U_{2,3}^{123},U_{1,1}^{4})$ $(\mathcal{W}_{2}^{34},U_{1,1}^{1})$ 44 44, Thm.IV-E.1 $(\hat{U}^{1}_{2,5},U_{1,1}^{2})$ $(\hat{U}^{1}_{3,5},U^{2}_{1,1})$ $(V^{12}_{8},U^{3}_{1,1})$
$12$, Thm.IV-A.1 44, Thm.IV-A.1 44, Thm.IV-B.2 $12$, Thm.IV-B.2 $12$, Thm.IV-B.3 $12$, Thm.IV-B.5 $12$, Thm.IV-G.1
$U^{\beta}_{1,2}$ $\backslash$ $(\mathcal{W}_{2}^{14},U_{1,2}^{14})$
$(U^{12}_{1,2},U^{13}_{1,2})$ $({U^{123}_{1,3},U^{12}_{1,2})}$ $(U_{2,3}^{123},U_{1,2}^{12})$ 66, Thm.IV-B.2
$(U^{12}_{1,2},U^{34}_{1,2})$ $({U^{123}_{1,3},U^{14}_{1,2})}$ 66, Thm.IV-A.1 $(U_{2,3}^{123},U_{1,2}^{14})$ $24$, Thm.IV-B.2 66 66, Thm.IV-E.2 $12$, Thm.IV-B.3 $12$, Thm.IV-B.5 $24$, Thm.IV-G.1
33, Thm.IV-A.1 $12$, Thm.IV-A.1 $12$, Thm.IV-B.2 $(\mathcal{W}_{2}^{34},U_{1,2}^{12})$
66, Thm.IV-B.1
@@
66, Thm.IV-A.1 44, Thm.IV-A.1 $12$, Thm.IV-C.1 $12$, Thm.IV-B.1 44 44, Thm.IV-E.4 $12$, Thm.IV-B.4 44, Thm.IV-F.1 $12$, Thm.IV-G.1
$U_{1,4}$ $\backslash \\backslash \\$backslash $(U_{2,3}^{123},U_{1,4})$ 0 0 $(U_{3,4},U_{1,4})$ 0 0 $(V^{12}_{8},U_{1,4})$
44, Thm.IV-C.3 11, Thm.IV-E.5 66, Thm.IV-G.1
$U^{\beta}_{2,3}$ $\backslash \\backslash \\$backslash $(U_{2,3}^{123},U_{2,3}^{124})$ $(\mathcal{W}_{2}^{12},U_{2,3}^{134})$ $(U_{2,4},U^{123}_{2,3}),4$ $(U_{3,4},U^{123}_{2,3})$ $(\hat{U}^{1}_{2,5},U_{2,3}^{234})$ $(\hat{U}^{1}_{3,5},U^{123}_{2,3})$ $(V^{12}_{8},U^{123}_{2,3})$
66, Thm.IV-C.2 $12$, Thm.IV-D.1 44 44, Thm.IV-E.3 44 $12$, Thm.IV-F.2 $12$, Thm.IV-G.1
$\mathcal{W}^{\beta}_{2}$ $\backslash \\backslash \\$backslash $\backslash$ $(\mathcal{W}_{2}^{12},\mathcal{W}_{2}^{13})$ $(U_{2,4},\mathcal{W}^{12}_{2})$ 0 $(\hat{U}^{1}_{2,5},\mathcal{W}_{2}^{12})$ $(\hat{U}^{1}_{3,5},\mathcal{W}^{23}_{2})$ 0
$12$ 66 $12$ $12$, Thm.IV-F.3
$U_{2,4}$ $\backslash \\backslash \\$$\backslash \\backslash \\$backslash 0 $(\hat{U}^{1}_{2,5},U_{2,4})$ $(\hat{U}^{1}_{3,5},U_{2,4})$ 0
44 44
$U_{3,4}$ $\backslash \\backslash \\$$\backslash \\backslash \\$backslash $\backslash$ 0 0 $(V^{12}_{8},U_{3,4})$
66, Thm.IV-G.1
$\hat{U}^{j}_{2,5}$ $\backslash \\backslash \\$$\backslash \\backslash \\$backslash $\backslash$ 0 0 0
$\hat{U}^{j}_{3,5}$ $\backslash \\backslash \\$$\backslash \\backslash \\$$\backslash \\backslash$ 0 0
$V^{\beta}_{8}$ $\backslash \\backslash \\$$\backslash \\backslash \\$$\backslash \\backslash \\$backslash 0
@@ IV Characterization of entropy functions on two-dimensional faces of $\Gamma_{4}$
In this section, we characterize the entropy functions on $2$-dimensional faces of $\Gamma_{4}$. We embed each face $F=(E_{1},E_{2})$ in the first octant of a $2$-dimensional cartesian coordinate system whose axis are labeled by $a$ and $b$. Thus, for each $(a,b)$, $a,b\geq 0$, it represents the polymatroid $a\mathbf{r}_{1}+b\mathbf{r}_{2}$, where $\mathbf{r}_{i},i=1,2$ is the rank function of the minimal integer polymatroid in $E_{i}$, respectively. Throughout this paper, for a random vector $(X_{i},i\in N_{4})$ or its subvectors, we assume each $X_{i}$ is distributed on a finite set $\mathcal{X}_{i}$, and for each $x_{i}\in$ $\mathcal{X}_{i}$, $p(x_{i})>0$.
@@
For $(x_{i},i\in N_{3})\in\mathcal{X}_{N_{3}}$ with $p(x_{123})>0$, the information equalities imply that the probability mass function satisfies
@@
Now we prove that each connected component is a complete tripartite graph. In a connected component, for any two vertices $x_{i}\in\mathcal{X}_{i},x_{j}\in\mathcal{X}_{j}$, $i,j\in N_{3},i\neq j$, there exists a path from $x_{i}$ to $x_{j}$ and the length of the path is denoted by $n$. We prove that $x_{i}$ and $x_{j}$ are adjacent by induction on $n$. First, it is obviously ture for $n=1$ and it holds for $n=2$ by Claim 1. Assume it is ture for some positive integers k1k-1 and $k$ where $k\geq 2$. Now consider $n=k+1$. If the vertex $y$ adjacent to $x_{j}$ on the path is not in $\mathcal{X}_{i}$, by induction hypothesis on $k$, we see that $x_{i}$ is adjacent to $y$. Then by Lemma 1, $x_{i}$ is adjacent to $x_{j}$. If $y$ is in $\mathcal{X}_{i}$, the vertex $z(z\neq x_{j})$ adjacent to $y$ on the path is adjacent to $x_{i}$ by induction on k1k-1. Then by Claim 2, $x_{i}$ is adjacent to $x_{j}$.
@@ IV-A All-entropic faces
Theorem IV-A.1 is also a corollary of [33, Proposition 4] or [34, Theorem 2].
@@ IV-B Entropy functions on the faces extended from $2$-dimensional faces of $\Gamma_{3}$
For $F=(U_{2,3}^{123},U_{1,2}^{12}),(\mathcal{W}_{2}^{34},U_{1,2}^{12})$, and $(\mathcal{W}_{2}^{14},U_{1,3}^{124})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a+b\geq\log\lceil 2^{a}\rceil$. [^4] We take the base of the logarithm be $2$ for computing the entropy.
@@
Let $F$ be any of the $3$ faces and $\mathbf{h}\in F$. Let $\mathbf{h}=a\mathbf{r}_{1}+b\mathbf{r}_{2}$, where $\mathbf{r}_{1}$ and $\mathbf{r}_{2}$ are the rank functions of the matroids on the two extreme rays of the face, respectively. Restricting $\mathbf{h}$ on $N_{3}$, we obtain $\mathbf{h}^{\prime}=a\mathbf{r}^{\prime}_{1}+b\mathbf{r}^{\prime}_{2}$, where $\mathbf{r}^{\prime}_{i},i=1,2$ are the restriction of $\mathbf{r}_{i}$ on $N_{3}$. It can be checked that they are the rank functions of $U_{2,3}$ and $U^{12,3}_{1,2}$, respectively. Thus, $\mathbf{h}^{\prime}\in(U_{2,3},U^{12,3}_{1,2})$ and $a+b\geq\log\lceil 2^{a}\rceil$ by [10, Theorem 1].
@@
For $F=(U_{2,3}^{123},U_{1,1}^{1})$, $(U_{2,3}^{123},U_{1,1}^{4})$, $(U_{2,3}^{123},U_{1,2}^{14})$, $(\mathcal{W}_{2}^{14},U_{1,1}^{1})$, $(\mathcal{W}_{2}^{34},U_{1,1}^{1})$, $(\mathcal{W}_{2}^{14},U_{1,2}^{14})$ and $(\mathcal{W}_{2}^{24},U_{1,2}^{14})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a=\log{v}$ for integer $v>0$.
[…]
@@
For $F=(U_{2,3}^{123},U_{1,1}^{4})$ and $\mathbf{h}\in F$, restricting $\mathbf{h}$ on $N_{3}$, we obtain $\mathbf{h}^{\prime}=a\mathbf{r}^{\prime}_{1}$, where $\mathbf{r}^{\prime}_{1}$ is the rank function of $U_{2,3}$, the matroid on the extreme ray of $\Gamma_{3}$. Thus, if $\mathbf{h}$ is entropic, then $a=\log{v}$ for some integer $v>0$. Now for any $\mathbf{h}\in F$ with $a=\log{v}$, let $\mathbf{h}^{\prime}$ be its restriction on $N_{3}$ and $(X_{1},X_{2},X_{3})$ be the characterizing random variable of $\mathbf{h}^{\prime}$. Let $X_{4}$ be an arbitrary random variable independent of $(X_{1},X_{2},X_{3})$ with $H(X_{4})=b$. Then it can be checked that $(X_{i},i\in N_{4})$ is the characterizing random variable of $\mathbf{h}$. Thus $\mathbf{h}$ is entropic.
For $F$ be any of the other $6$ faces and $\mathbf{h}\in F$, let $\mathbf{h}=a\mathbf{r}_{1}+b\mathbf{r}_{2}$, where $\mathbf{r}_{1}$ and $\mathbf{r}_{2}$ are the rank functions of the matroids on the two extreme rays of the face, respectively. Restricting $\mathbf{h}$ on $N_{3}$, we obtain $\mathbf{h}^{\prime}=a\mathbf{r}^{\prime}_{1}+b\mathbf{r}^{\prime}_{2}$, where $\mathbf{r}^{\prime}_{i}=1,2$ is the restriction of $\mathbf{r}_{i}$ on $N_{3}$, and the rank function of $U_{2,3}$ and $U^{1,3}_{1,1}$, respectively. Thus, $\mathbf{h}^{\prime}\in(U_{2,3},U^{1,3}_{1,1})$ and $a=\log{v}$ for positive $v\in\mathbb{Z}$ by [11, Theorem 1].
Now we prove the “if ” part. Let $\mathbf{h}=a\mathbf{r}_{1}+b\mathbf{r}_{2}\in F$ and $a=\log{v}$ for some integer $v\geq 0$. Let $\mathbf{h}^{\prime}$ be the restriction of $\mathbf{h}$ on $N_{3}$. Therefore, $\mathbf{h}^{\prime}\in(U_{2,3},U^{1,3}_{1,1})$. Note that $\mathbf{h}^{\prime}=a\mathbf{r}_{1}+b\mathbf{r}_{2}$, where $\mathbf{r}_{1}$ and $\mathbf{r}_{2}$ are respectively the rank functions of $U_{2,3}$ and $U^{1,3}_{1,2}$. Let $(X^{\prime}_{1},X^{\prime}_{2},X^{\prime}_{3})$ be the characterizing random vector of $a\mathbf{r}_{1}$, that is, $X^{\prime}_{1}$ and $X^{\prime}_{2}$ are independent and both uniformly distributed on $\mathbb{I}_{v}$, and $X^{\prime}_{3}=X^{\prime}_{1}+X^{\prime}_{2}\mod v$. Let $(X^{\prime\prime}_{1},X^{\prime\prime}_{2},X^{\prime\prime}_{3})$ be characterizing random vector of $b\mathbf{r}_{2}$, that is, $X^{\prime\prime}_{1}$ is an arbitrary random variable with $H(X^{\prime\prime}_{1})=b$, and $X^{\prime\prime}_{2}$ and $X^{\prime\prime}_{3}$ are both constant. Let $X_{i}=(X^{\prime}_{i},X^{\prime\prime}_{i})$, $i\in N_{3}$. Then it can be checked that $(X_{1},X_{2},X_{3})$ is the characterizing random vector of $\mathbf{h}^{\prime}$.
@@
For $F=(\hat{U}_{2,5}^{1},U_{1,1}^{1})$, $(\hat{U}_{2,5}^{1},U_{1,1}^{2})$ and $(\hat{U}_{2,5}^{1},U_{1,2}^{12})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a=\log v$ for integer $v>0$.